# On Bits and Bandits: Quantifying the Regret-Information Trade-off

### v3

@@ 4.2 Entropy Constraint
While Proposition 4.1 shows how we can lower bound regret under a constraint on accumulated information, we now shift our focus to a constraint on the prior only, i.e $H(\pi^{*})\leq R$. Russo and Van Roy [2014, 2016] showed that in this scenario, the regret of Thompson sampling and information-directed sampling (IDS) can be upper bounded using $H(\pi^{*})$.
Theorem 4.4 (Propositions 2 and 4 of Russo and Van Roy [2014]).
@@
$D\left(p_{t},P^{\pi^{*}}\right)\geq c\bar{A}R$ for all $t$, for some constant $c>0$.
@@ 5 Experiments
5.1 Stochastic Bayesian Bandit
We begin by demonstrating the trade-off in a simple Bayesian MAB problem. In this problem, our model space contains $K$ possible models. The decision space of the problem also contains $K$ decisions, which are called arms. The mean of all arms in every model is $0.5(1-\varepsilon)$, except for one arm for which the mean is $0.5(1+\varepsilon)$. The optimal arm is different for every model, and the reward of each arm is Bernoulli distributed. We also fix $K=8,\varepsilon=0.1$ and a uniform prior over the problems and compare three different bandit algorithms - EXP3 [Lattimore and Szepesvári, 2020], APS [Xu and Zeevi, 2023], and Thompson sampling [Thompson, 1933]. We also estimate the accumulated information for each algorithm. Results are presented in Figure 2. We see how algorithms that accumulate information quickly also suffer less regret. We also compare Thompson sampling under different priors $H(\pi^{*})$. The results of this experiment can be found in Figure 3. Additional experimental details are provided in Appendix E.
@@ 5.2 Question Answering with Large Language Models
We run the experiment described above with the following specifications. To generate the multiple-choice questions, we used the MMLU dataset, which contains multiple-choice questions in a variety of topics, such as algebra, business, etc. [Hendrycks et al., 2021a, b].
@@
% Deployment Bits-based Random
Mistral 7B $50\%$ $\mathbf{36\pm 1}$ $38.7\pm 0.3$
@@
We used 10 different seeds to generate 10 sets of 200 questions from MMLU randomly. For the small LLM we either used Mistral 7B Jiang et al. [2023], Falcon 7B [Almazrouei et al., 2023], or Gemma 2B [Team et al., 2024]. We used Mixtral 7Bx8 [Jiang et al., 2024], Llama3-70B [AI@Meta, 2024], or Gemma 7B for the large LLM. We applied 4-bit quantization [Dettmers and Zettlemoyer, 2023] for all models and flash attention [Dao et al., 2022] for all models excluding Falcon 7B. Table 2 describes the mean regret of every policy under a different number of large LLM deployments. Tables 4 and 5 provide additional results for different combinations of small and large LLMs.
@@ 6 Related Work
In the Bayesian setting, prior information is assumed to be expressed as a probability distribution, called the prior. New information is then incorporated using Bayesian inference to update the prior probability, resulting in a posterior distribution. Bayesian algorithms have been extensively studied in multiple online decision-making problems such as multi-armed bandits [Russo and Van Roy, 2014, Kaufmann et al., 2012] and reinforcement learning [Guez et al., 2012, Ghavamzadeh et al., 2016]. Upper bounds for Bayesian algorithms have been proved for both the frequentist [Kaufmann et al., 2011] and Bayesian settings [Russo and Van Roy, 2014, 2016]. Prior dependent regret bounds for Bayesian algorithms have also been studied [Bubeck and Liu, 2013, Russo and Van Roy, 2014]. Previous studies have explored prior-dependent lower bounds for Bayesian regret, but these are limited to specific forms of priors, such as Gaussian priors [Atsidakou et al., 2024].
@@
In the contextual setting, information is revealed to the agent before every round, which can be leveraged to minimize regret. This framework has gathered attention within both the bandit [Bietti et al., 2021] and reinforcement learning frameworks [Klink et al., 2020, Modi and Tewari, 2020]. The contexts can be selected arbitrarily by an adversary [Beygelzimer et al., 2011, Chu et al., 2011] or generated from some prior probability distribution, similarly to the Bayesian setting [Hao et al., 2020, May et al., 2012]. Bayesian algorithms have also been adapted to the arbitrary context setting [Agrawal and Goyal, 2013]. Furthermore, the type of information provided to the learner can expand beyond the contextual information provided directly [Schneider and Zimmert, 2024]. Our work introduces a contextual Bayesian setting framework that covers a wide variety of interactive decision-making tasks.
@@
have been employed to derive upper bounds for various online tasks. Russo and Van Roy [2016] introduced an information-theoretic method for establishing Bayesian regret upper bound for Thompson sampling in the multi-armed bandit setting, which depends on the entropy of the prior. This approach was also used to design new Bayesian algorithms that utilize mutual information called information-directed sampling (IDS) [Russo and Van Roy, 2014]. IDS and the information-theoretic method it utilizes were also extended to other tasks such as contextual bandits [Neu et al., 2022], sparse linear bandits [Hao et al., 2021], non-stationary bandits [Min and Russo, 2023, Liu et al., 2023], reinforcement learning [Lu and Van Roy, 2019], non-linear control [Kakade et al., 2020], partial monitoring [Lattimore and Szepesvári, 2019], and other online optimization problems [Liu et al., 2018, Dong et al., 2019, Lattimore and Gyorgy, 2021, Russo and Van Roy, 2018]. The bounds presented in works such as [Russo and Van Roy, 2016, 2014, Neu et al., 2022] are based on the upper bounding of the information ratio. Our work uses a different approach, which obtains nearly matching prior-dependent lower bounds for Thompson sampling and information-directed sampling (IDS). Furthermore, these bounds can be applied beyond the bandit setting. Seldin et al. [2011] utilized PAC-Bayes bounds to obtain information-theoretic upper bounds on the per-round regret, that depend on mutual information. Arumugam and Van Roy [2021, 2022] have explored a different information-theoretic method, utilizing rate-distortion theory to minimize regret.
@@ 7 Conclusions and Future Work
A central assumption in our analysis is that exogenous information can be gathered regardless of the process history. However, this assumption may be violated in a general sequential decision process where the action itself may have ramifications on the quality, quantity, and cost of the exogenous information. Our work only covers the Bayesian setting, and our definition of the measure of information relies on it. In other settings, such as adversarial learning [Neu and Olkhovskaya, 2020] or frequentist settings, a different measure of information is required. Another limitation is the difference in $\sqrt{\log K}$ between the lower and upper bounds presented in this work. Making these bounds tighter can be the topic of future work.
@@ Appendix A Table of Notations
$p_{t}$ Policy selected at round $t$.
@@ Appendix B Useful Properties of Covering Numbers
Definition B.1 (Covering and packing numbers [Wainwright, 2019]).
@@
  • (Theorem 14.2 of Wu [2020]) For any subset of dimension $d$ that is contained in $\mathbb{B}_{p}^{d}$ the covering number $\mathcal{N}_{p}(\epsilon)$ obeys $d\log\left(\frac{2}{\epsilon}\right)\leq\log\mathcal{N}_{p}(\epsilon)\leq d\log\left(\frac{3}{\epsilon}\right)$. This includes the probability simplex with $d$ variables and $\mathbb{B}_{d}^{2}$.
  • (Theorem 14.1 of Wu [2020]) Under the same norms, for the same set,
@@ C.3 Proofs for Additional Lower Bounds
We focus on the set $\Delta(\Pi)$ and note that it is a parametric set. We also see that for all policies $\phi,\psi\in\Delta(\Pi)$ we have that
@@
Theorem C.4 (Theorem 1 of Lattimore and Szepesvári [2019]).
@@ C.4 Additional Explanation
This decision stems from the improved upper bound of $I(V;\mathcal{H}_{T})$ for parametric decision space, which is found in Theorem 3.3. Using the other upper bound for $I(V;\mathcal{H}_{T})$ simply results in a sub-optimal lower bound. Additional details can be found in Yang and Barron [1999].
@@ Appendix D Proofs for Section 4
The proof follows a similar analysis to the proof of Theorem 3.3, which is provided in Yang and Barron [1999]. Since we focus on a parametric set, we consider a local packing set with a cardinality of $\mathcal{M}^{\mathrm{loc}}(\rho,\epsilon,\Delta(\Pi))$, which we denote by $E$. $\tilde{E}$ is the set contained by $E$ under the mutual information constraint,
@@
$=\sum_{t=1}^{T}D\left(P^{V\mid\mathcal{H}_{T}},P^{V\mid\mathcal{H}_{t}}\right)$
$\leq T\max_{P_{1},P_{2}\in\tilde{E}}D\left(P_{1},P_{2}\right)$
$\leq T\frac{R}{\log K}\max_{P_{1},P_{2}\in E}D\left(P_{1},P_{2}\right)$
@@
We consider the following two algorithms. The first is Thompson sampling, whose policy at round $t$ is the posterior probability, $p_{t}=P^{\pi^{*}\mid\mathcal{H_{t}}}$. The second algorithm is Thompson sampling with the constrained observations $\tilde{O}_{t}$ such that $I(\pi^{*};\tilde{\mathcal{H}}_{T})=D\left(P^{\pi^{*}\mid\tilde{\mathcal{H}}_{T}},P^{\pi^{*}}\right)=R$, where ~$\tilde{\mathcal{H}}_{t}$ is the history with constrained observations. Now, we define the information ratio for both the normal and constrained variants of Thompson sampling:
@@
$\max_{P}\max_{t}\tilde{\rho}_{t}$ $\leq\max_{P}\max_{t}\frac{\log K}{D\left(P^{\pi^{*}\mid(\tilde{O_{t}},\tilde{\mathcal{H}}_{t})},P^{\pi^{*}\mid\tilde{\mathcal{H}}_{t}}\right)}\rho_{t}$
@@
Where the third inequality follows by the information ratio upper bound presented by Russo and Van Roy [2016]. If the MAB is linear then the information ratio $\rho_{t}$ can be upper-bounded by $\frac{d}{2}$. Then, we have the following upper bound,
@@
Now, by the analysis done by Russo and Van Roy [2016], we know that if $\max_{t}\tilde{\rho}_{t}\leq\tilde{\rho}$ for all $t$, then the Bayesian regret can be upper bounded by
@@
Let $P_{1},P_{2}$ be two priors such that $H(P_{1}^{\pi^{*}})=R\leq\log K$. Also let $\psi,\phi$ be two decisions such that $D\left(\psi,P_{1}^{\pi^{*}}\right)\geq c\bar{A}R$ and $D\left(\phi,P_{2}^{\pi^{*}}\right)\leq\log K$. Then,
@@
$(\mathbb{E}_{P_{2}}[\mu_{M}(\pi^{*},C)-\mu_{M}(\phi,C)])^{2}\leq 2D\left(\phi,P_{2}^{\pi^{*}}\right)\leq\log K.$
@@
for any prior $\nu$ and policies $p^{\prime}_{t}$ such that $D\left(p^{\prime}_{t},\nu^{\pi^{*}}\right)\leq\log K$. So we can take the worst-case regret,
$\mathcal{BR}_{P}(T;{p_{t}})$ $\geq 2\sqrt{c\frac{R}{\log K}}\max_{\nu}\inf_{\{p^{\prime}_{t}\}:D\left(p^{\prime}_{t},\nu^{\pi^{*}}\right)\leq\log K}\sum_{t=1}^{T}\mathbb{E}_{\nu}[\mu_{M}(\pi^{*},C_{t})-\mu_{M}(p_{t}^{\prime},C_{t})]$
@@
where the second inequality follows by removing a constraint on the available policies. ∎
@@
Let there be a $K$-MAB. Then for any algorithm such that $D\left(p_{t},P^{\pi^{*}}\right)\geq c\bar{A}H(\pi^{*})$, the regret can be lower-bounded by $\Omega\left(\sqrt{\frac{KH(\pi^{*})T}{\log K}}\right)$. Additionally, if the MAB is linear, then the regret can be lower bounded by $\Omega\left(\sqrt{\frac{dH(\pi^{*})T}{\log K}}\right)$.
@@
$D\left(p_{t},P^{\pi^{*}}\right)\geq(1-\delta)H(\pi^{*})-c_{\delta}$
@@
$D\left(p_{t},P^{\pi^{*}}\right)$ $=\mathbb{E}\left[\sum_{\pi}p_{t}(\pi)\log\left(\frac{p_{t}(\pi)}{P^{\pi^{*}}(\pi)}\right)\right]$
@@
From this, we can conclude that any algorithm whose policy converges to the optimal decision obeys the assumption.

### v4

@@ 4.2 Entropy Constraint
While Proposition 4.1 shows how we can lower bound regret under a constraint on accumulated information, we now shift our focus to a constraint on the prior only, i.e $H(\pi^{*})\leq R$. Russo & Van Roy (2014; 2016) showed that in this scenario, the regret of Thompson sampling and information-directed sampling (IDS) can be upper bounded using $H(\pi^{*})$.
Theorem 4.4 (Propositions 2 and 4 of Russo & Van Roy (2014)).
@@
$D_{\mathrm{KL}}\left(p_{t}\parallel P^{\pi^{*}}\right)\geq c\bar{A}R$ for all $t$, for some constant $c>0$.
@@ 5 Experiments
5.1 Stochastic Bayesian Bandit
We begin by demonstrating the trade-off in a simple Bayesian MAB problem. In this problem, our model space contains $K$ possible models. The decision space of the problem also contains $K$ decisions, which are called arms. The mean of all arms in every model is $0.5(1-\varepsilon)$, except for one arm for which the mean is $0.5(1+\varepsilon)$. The optimal arm is different for every model, and the reward of each arm is Bernoulli distributed. We also fix $K=8,\varepsilon=0.1$ and a uniform prior over the problems and compare three different bandit algorithms - EXP3 (Lattimore & Szepesvári, 2020), APS (Xu & Zeevi, 2023), and Thompson sampling (Thompson, 1933). We also estimate the accumulated information for each algorithm. Results are presented in Figure 2. We see how algorithms that accumulate information quickly also suffer less regret. We also compare Thompson sampling under different priors $H(\pi^{*})$. The results of this experiment can be found in Figure 3. Additional experimental details are provided in Appendix E.
@@ 5.2 Question Answering with Large Language Models
We run the experiment described above with the following specifications. To generate the multiple-choice questions, we used the MMLU dataset, which contains multiple-choice questions in a variety of topics, such as algebra, business, etc. (Hendrycks et al., 2021b; a).
@@
% Deployment Bits-based Random Small only Large only
Mistral 7B $50\%$ $\mathbf{36\pm 1}$ $38.7\pm 0.3$ $59.3\pm 0.7$ $18.2\pm 0.6$
@@
We used 10 different seeds to generate 10 sets of 200 questions from MMLU randomly. For the small LLM we either used Mistral 7B Jiang et al. (2023), Falcon 7B (Almazrouei et al., 2023), or Gemma 2B (Team et al., 2024). We used Mixtral 7Bx8 (Jiang et al., 2024), Llama3-70B (AI@Meta, 2024), or Gemma 7B for the large LLM. We applied 4-bit quantization (Dettmers & Zettlemoyer, 2023) for all models and flash attention (Dao et al., 2022) for all models excluding Falcon 7B. Table 2 describes the mean regret of every policy under a different number of large LLM deployments. Tables 4 and 5 provide additional results for different combinations of small and large LLMs.
@@ 6 Related Work
In the Bayesian setting, prior information is assumed to be expressed as a probability distribution, called the prior. New information is then incorporated using Bayesian inference to update the prior probability, resulting in a posterior distribution. Bayesian algorithms have been extensively studied in multiple online decision-making problems such as multi-armed bandits (Russo & Van Roy, 2014; Kaufmann et al., 2012) and reinforcement learning (Guez et al., 2012; Ghavamzadeh et al., 2016). Upper bounds for Bayesian algorithms have been proved for both the frequentist (Kaufmann et al., 2011) and Bayesian settings (Russo & Van Roy, 2014; 2016). Prior dependent regret bounds for Bayesian algorithms have also been studied (Bubeck & Liu, 2013; Russo & Van Roy, 2014). Previous studies have explored prior-dependent lower bounds for Bayesian regret, but these are limited to specific forms of priors, such as Gaussian priors (Atsidakou et al., 2024).
@@
In the contextual setting, information is revealed to the agent before every round, which can be leveraged to minimize regret. This framework has gathered attention within both the bandit (Bietti et al., 2021) and reinforcement learning frameworks (Klink et al., 2020; Modi & Tewari, 2020). The contexts can be selected arbitrarily by an adversary (Beygelzimer et al., 2011; Chu et al., 2011) or generated from some prior probability distribution, similarly to the Bayesian setting (Hao et al., 2020; May et al., 2012). Bayesian algorithms have also been adapted to the arbitrary context setting (Agrawal & Goyal, 2013). Furthermore, the type of information provided to the learner can expand beyond the contextual information provided directly (Schneider & Zimmert, 2024). Our work introduces a contextual Bayesian setting framework that covers a wide variety of interactive decision-making tasks.
@@
have been employed to derive upper bounds for various online tasks. Russo & Van Roy (2016) introduced an information-theoretic method for establishing Bayesian regret upper bound for Thompson sampling in the multi-armed bandit setting, which depends on the entropy of the prior. This approach was also used to design new Bayesian algorithms that utilize mutual information called information-directed sampling (IDS) (Russo & Van Roy, 2014). IDS and the information-theoretic method it utilizes were also extended to other tasks such as contextual bandits (Neu et al., 2022), sparse linear bandits (Hao et al., 2021), non-stationary bandits (Min & Russo, 2023; Liu et al., 2023), reinforcement learning (Lu & Van Roy, 2019), non-linear control (Kakade et al., 2020), partial monitoring (Lattimore & Szepesvári, 2019), and other online optimization problems (Liu et al., 2018; Dong et al., 2019; Lattimore & Gyorgy, 2021; Russo & Van Roy, 2018). The bounds presented in works such as (Russo & Van Roy, 2016; 2014; Neu et al., 2022) are based on the upper bounding of the information ratio. Our work uses a different approach, which obtains nearly matching prior-dependent lower bounds for Thompson sampling and information-directed sampling (IDS). Furthermore, these bounds can be applied beyond the bandit setting. Seldin et al. (2011) utilized PAC-Bayes bounds to obtain information-theoretic upper bounds on the per-round regret, that depend on mutual information. Arumugam & Van Roy (2021; 2022) have explored a different information-theoretic method, utilizing rate-distortion theory to minimize regret.
@@ 7 Conclusions and Future Work
A central assumption in our analysis is that exogenous information can be gathered regardless of the process history. However, this assumption may be violated in a general sequential decision process where the action itself may have ramifications on the quality, quantity, and cost of the exogenous information. Our work only covers the Bayesian setting, and our definition of the measure of information relies on it. In other settings, such as adversarial learning (Neu & Olkhovskaya, 2020) or frequentist settings, a different measure of information is required. Another limitation is the difference in $\sqrt{\log K}$ between the lower and upper bounds presented in this work. Making these bounds tighter can be the topic of future work.
@@ Appendix A Table of Notations
$p_{t}$ stochastic decision selected at round $t$.
@@ Appendix B Useful Properties of Covering Numbers
Definition B.1 (Covering and packing numbers (Wainwright, 2019)).
@@
  • (Theorem 14.2 of Wu (2020)) For any subset of dimension $d$ that is contained in $\mathbb{B}_{p}^{d}$ the covering number $\mathcal{N}_{p}(\epsilon)$ obeys $d\log\left(\frac{2}{\epsilon}\right)\leq\log\mathcal{N}_{p}(\epsilon)\leq d\log\left(\frac{3}{\epsilon}\right)$. This includes the probability simplex with $d$ variables and $\mathbb{B}_{d}^{2}$.
  • (Theorem 14.1 of Wu (2020)) Under the same norms, for the same set,
@@ C.3 Proofs for Additional Lower Bounds
We focus on the set $\Delta(\Pi)$ and note that it is a parametric set. We also see that for all stochastic decisions $\phi,\psi\in\Delta(\Pi)$ we have that
@@
Theorem C.4 (Theorem 1 of Lattimore & Szepesvári (2019)).
@@ C.4 Additional Explanation
This decision stems from the improved upper bound of $I(V;\mathcal{H}_{T})$ for parametric decision space, which is found in Theorem 3.3. Using the other upper bound for $I(V;\mathcal{H}_{T})$ simply results in a sub-optimal lower bound. Additional details can be found in Yang & Barron (1999).
@@ Appendix D Proofs for Section 4
The proof follows a similar analysis to the proof of Theorem 3.3, which is provided in Yang & Barron (1999). Since we focus on a parametric set, we consider a local packing set with a cardinality of $\mathcal{M}^{\mathrm{loc}}(\rho,\epsilon,\Delta(\Pi))$, which we denote by $E$. $\tilde{E}$ is the set contained by $E$ under the mutual information constraint,
@@
$=\sum_{t=1}^{T}D_{\mathrm{KL}}\left(P^{V\mid\mathcal{H}_{T}}\parallel P^{V\mid\mathcal{H}_{t}}\right)$
$\leq T\max_{P_{1},P_{2}\in\tilde{E}}D_{\mathrm{KL}}\left(P_{1}\parallel P_{2}\right)$
$\leq T\frac{R}{\log K}\max_{P_{1},P_{2}\in E}D_{\mathrm{KL}}\left(P_{1}\parallel P_{2}\right)$
@@
We consider the following two algorithms. The first is Thompson sampling, which means that, $p_{t}=P^{\pi^{*}\mid\mathcal{H_{t}}}$. The second algorithm is Thompson sampling with the constrained observations $\tilde{O}_{t}$ such that $I(\pi^{*};\tilde{\mathcal{H}}_{T})=D_{\mathrm{KL}}\left(P^{\pi^{*}\mid\tilde{\mathcal{H}}_{T}}\parallel P^{\pi^{*}}\right)=R$, where ~$\tilde{\mathcal{H}}_{t}$ is the history with constrained observations. Now, we define the information ratio for both the normal and constrained variants of Thompson sampling:
@@
$\max_{P}\max_{t}\tilde{\rho}_{t}$ $\leq\max_{P}\max_{t}\frac{\log K}{D_{\mathrm{KL}}\left(P^{\pi^{*}\mid(\tilde{O_{t}},\tilde{\mathcal{H}}_{t})}\parallel P^{\pi^{*}\mid\tilde{\mathcal{H}}_{t}}\right)}\rho_{t}$
@@
Where the third inequality follows by the information ratio upper bound presented by Russo & Van Roy (2016). If the MAB is linear then the information ratio $\rho_{t}$ can be upper-bounded by $\frac{d}{2}$. Then, we have the following upper bound,
@@
Now, by the analysis done by Russo & Van Roy (2016), we know that if $\max_{t}\tilde{\rho}_{t}\leq\tilde{\rho}$ for all $t$, then the Bayesian regret can be upper bounded by
@@
Let $P_{1},P_{2}$ be two priors such that $H(P_{1}^{\pi^{*}})=R\leq\log K$. Also let $\psi,\phi$ be two decisions such that $D_{\mathrm{KL}}\left(\psi\parallel P_{1}^{\pi^{*}}\right)\geq c\bar{A}R$ and $D_{\mathrm{KL}}\left(\phi\parallel P_{2}^{\pi^{*}}\right)\leq\log K$. Then,
@@
$(\mathbb{E}_{P_{2}}[\mu_{M}(\pi^{*},C)-\mu_{M}(\phi,C)])^{2}\leq 2D_{\mathrm{KL}}\left(\phi\parallel P_{2}^{\pi^{*}}\right)\leq\log K.$
@@
for any prior $\nu$ and stochastic decisions $p^{\prime}_{t}$ such that $D_{\mathrm{KL}}\left(p^{\prime}_{t}\parallel\nu^{\pi^{*}}\right)\leq\log K$. So we can take the worst-case regret,
$\mathcal{BR}_{P}(T;{p_{t}})$ $\geq 2\sqrt{c\frac{R}{\log K}}\max_{\nu}\inf_{\{p^{\prime}_{t}\}:D_{\mathrm{KL}}\left(p^{\prime}_{t}\parallel\nu^{\pi^{*}}\right)\leq\log K}\sum_{t=1}^{T}\mathbb{E}_{\nu}[\mu_{M}(\pi^{*},C_{t})-\mu_{M}(p_{t}^{\prime},C_{t})]$
@@
where the second inequality follows by removing a constraint on the available decisions. ∎
@@
Let there be a $K$-MAB. Then for any algorithm such that $D_{\mathrm{KL}}\left(p_{t}\parallel P^{\pi^{*}}\right)\geq c\bar{A}H(\pi^{*})$, the regret can be lower-bounded by $\Omega\left(\sqrt{\frac{KH(\pi^{*})T}{\log K}}\right)$. Additionally, if the MAB is linear, then the regret can be lower bounded by $\Omega\left(\sqrt{\frac{dH(\pi^{*})T}{\log K}}\right)$.
@@
$D_{\mathrm{KL}}\left(p_{t}\parallel P^{\pi^{*}}\right)\geq(1-\delta)H(\pi^{*})-c_{\delta}$
@@
$D_{\mathrm{KL}}\left(p_{t}\parallel P^{\pi^{*}}\right)$ $=\mathbb{E}\left[\sum_{\pi}p_{t}(\pi)\log\left(\frac{p_{t}(\pi)}{P^{\pi^{*}}(\pi)}\right)\right]$
@@
From this, we can conclude that any algorithm whose stochastic decision at round $t$ converges to the optimal decision obeys the assumption.